OpenStudy (calculator):

Integrate 1/(1+sqrt(x))

OpenStudy (calculator):

\[\Huge \int\limits \frac{1}{1+\sqrt{x}}dx\]

OpenStudy (calculator):

not so easy this one

OpenStudy (anonymous):

what is this.... calculus???

OpenStudy (calculator):

yes

OpenStudy (calculator):

@Hero @experimentX

OpenStudy (anonymous):

ohhh -.- imin algebra 1 can u help me?

hartnn (hartnn):

i would do this \(\Huge \int\limits \frac{1}{1+\sqrt{x}}dx=\Huge \int\limits \dfrac{\sqrt x}{\sqrt x} \frac{1}{1+\sqrt{x}}dx \\ \Huge= \int\limits \frac{\sqrt x+1-1}{\sqrt x(1+\sqrt{x})}dx\)

OpenStudy (calculator):

okay, i'm following

hartnn (hartnn):

\( \Huge= \int\limits \frac{\sqrt x+1-1}{\sqrt x(1+\sqrt{x})}dx \\ \huge =\int \dfrac{dx}{\sqrt x}-\int \dfrac{dx}{\sqrt x(1+\sqrt x)} \) now you can put u= \(\sqrt x+1\) in 2nd integral

hartnn (hartnn):

did u get that ?

OpenStudy (calculator):

yes

hartnn (hartnn):

could you now integrate both the integrals ?

OpenStudy (experimentx):

looks like you can use trigs also